### v1

# Primes and polygonal numbers

Abstract.
For all integers $r,s>2$, there are infinitely many primes which can be written as the sum of an $r$-gonal number and an $s$-gonal number. In general, any linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes. However, unless $r=s=4$, such primes can not be described by finitely many congruences.
Key words and phrases:
prime, polygonal number, triangular number, arithmetic progression
2020 Mathematics Subject Classification:
Primary 11N32, 11N13, 11N36; Secondary 11P32, 11B25, 11B05
1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]). Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares. It is natural to wonder whether the same holds for triangular numbers or more generally, whether for all integers $r,s\geq 3$, there are infinitely many primes which can be represented as the sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see Chapter 2 in [2])
$T(r,m):=(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
Even more generally, one may ask whether for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, there are infinitely many primes of the form
$\mu T(r,m)+\nu T(s,n).$
We answer this in affirmative.
Theorem 1.
For sufficiently large $N$, for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, the number of primes of the form
$\mu T(r,m)+\nu T(s,n)$
which are less than $N$ is
$\asymp\begin{cases}\dfrac{N}{\log N},&\text{ if }r=s=4\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{otherwise.}\end{cases}$
By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
Corollary 1.
Unless $r=s=4$, for each pair of mutually coprime positive integers $\mu$ and $\nu$, the primes of the form
$\mu T(r,m)+\nu T(s,n),$
can not be described as the primes contained in a set of admissible arithmetic progressions.
In other words, one can only describe the primes of the form
$ax^{2}+by^{2}$
by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply that for any pair of $a,b\in\mathbb{N}$, if we replace either $x^{2}$ or $y^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers misses infinitely many primes in every admissible arithmetic progression.
In particular, the sequence $\{4k+1\}_{k\in\mathbb{N}}$ or its subprogressions are the only admissible arithmetic progressions such that each prime in those progressions is a sum of two polygonal numbers.
It follows from Theorem 1 and Chebyshev’s theorem (Corollary 2.6 in [7]) that the set of the primes, which are sum of two polygonal numbers (or more generally, linear combination of two polygonal numbers with coprime positive integer coefficients) at least one of which is not a square, does not have positive relative upper density with respect to the set of all primes. So, Green-Tao Theorem [4] which guarantees the existence of arbitrarily long arithmetic progressions among the subsets of the primes with positive relative upper densities, is not applicable here either.
From Theorem 1, for $\mu=\nu=1$ and $r=s=3$, we obtain the following corollary which answers the Open Problem 1 in [1].
Corollary 2.
There are infinitely many primes which are sum of two triangular numbers.
Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]), we have
Corollary 3.
There are infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares.
It is natural to wonder whether there are infinitely many primes $p$ such that both $p$ and $14p+$ 1 are sums of two squares or more generally, whether both the residue classes $1$ and $3$ modulo $4$ contain infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares. Even more generally, one may ask:
Open Question 1.
Does every admissible arithmatic progression contain infinitely many primes which are sum of two triangular numbers?
2. Proof of the main result
The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below.
Theorem A (Theorem 1 in [5]).
A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial of degree $2$ in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\gg\frac{N}{\log N},$ (2.1)
if $\Delta:=b^{2}-4ac$ is a perfect square or $D:=af^{2}-bef+ce^{2}+g\Delta=0$, and
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\asymp\frac{N}{(\log N)^{3/2}},$ (2.2)
if $\Delta$ is not a perfect square and $D\neq 0$.
Lemma B (Lemma 1 in [5]).
B Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g\in\mathbb{Z}[x,y]$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
$\Delta=b^{2}-4ac=0,\ \alpha:=bf-2ce=0\text{ \ and \ }\beta:=be-2af=0.$
Proof of Theorem 1: For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
$\mu T(r,m)+\nu T(s,n),$
we note that (see (1.1)) the above linear combination is equal to
$P_{1}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu(4-s)y$ (2.3)
for $m=2x$ and $n=2y$,
$P_{2}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu sy+\nu$ (2.4)
for $m=2x$ and $n=2y+1$,
$P_{3}(x,y):=2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu rx+\nu sy+\mu+\nu$ (2.5)
for $m=2x+1$ and $n=2y+1$.
Since the coefficients of $x^{2}$ and $y^{2}$ in each of the above polynomials are positive and since the coefficient of $xy$ is zero, it follows that
$\Delta<0$ (2.6)
for each of the above polynomials (see Lemma B). Hence, $\frac{\partial{P_{i}}}{\partial x}$ and $\frac{\partial{P_{i}}}{\partial y}$ linearly independent for all $i\in\{1,2,3\}$.
Also, from (2.3), (2.4) and (2.5), we note that for each of the polynomials $P_{1},P_{2}$ and $P_{3}$, we have
$D=2\mu\nu^{2}(r-2)(4-s)^{2}+2\mu^{2}\nu(s-2)(4-r)^{2},$ (2.7)
which equals zero if and only if $r=s=4$ (see Theorem A).
In order to apply Theorem A, we must also check whether the polynomials $P_{1},P_{2}$ $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$. Suppose, for some $i\in\{1,2,3\}$, there exists $a,b,c,a^{\prime},b^{\prime},c^{\prime}\in\mathbb{Z}$ such that
$P_{i}(x,y)=(ax+by+c)(a^{\prime}x+b^{\prime}y+c^{\prime}).$
Comparing the coefficients of $x^{2}$, $y^{2}$ and $xy$ in both of the above expressions for $P_{i}(x,y)$, we obtain
$aa^{\prime}=2\mu(r-2),\ bb^{\prime}=2\nu(s-2)\text{ and }ab^{\prime}+a^{\prime}b=0.$
Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$. Hence, substituting $kb$ for $a$ and $-kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
$k^{2}=-\frac{\mu(r-2)}{\nu(s-2)}<0,$
which is absurd! Thus, $P_{i}(x,y)$ is irreducible in $\mathbb{Q}[x,y]$ for all $i\in\{1,2,3\}$. Hence, it follows from Gauss’s Lemma (see Ch.VI, §2 in [6]) that $P_{1},P_{2}$ and $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$ if and only if these are primitive, i.e. the greatest common divisor of the coefficients of each of these polynomials is 1.
Since for all $i\in\{1,2,3\}$, the coefficients of $x^{2}$ and $y^{2}$ in $P_{i}(x,y)$ are positive and even, whereas the coefficient of $xy$ is zero, it follows that if any of the remaining coefficients of $P_{i}$ are odd, then $P_{i}(x,y)$ produces infinitely many odd numbers for $x,y\in\mathbb{N}$. In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers. In the following we compute the gcd $G$ of the coefficients of $P_{i}$ for $i\in\{1,2,3\}$.
Case I. For $P_{1}$, from (2.3) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu(4-s))=\gcd(4,\mu r,\nu s).$
Hence, $G=1$ if and only if either $\mu r$ or $\nu s$ is odd. In particular, $G=1$ implies $D>0$ (see (2.7)). So, it follows from Theorem A that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{1}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Case II. For $P_{2}$, from (2.4) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu s,\nu)=\gcd(4,r,\nu).$
Hence, $G=1$ if and only if either $r$ or $\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Case III. For $P_{3}$, from (2.5) we note that
$G=\gcd(2\mu(r-2),2\nu(s-2),\mu r,\nu s,\mu+\nu)=\gcd(4,r,s,\mu+\nu).$
Hence, $G=1$ if and only if either $r$ or $s$ of $\mu+\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
Since either $\nu$ or $\mu+\nu$ is odd for all mutually coprime pairs $\mu$ and $\nu$, the claim follows from the above cases. ∎