### v4

# Primes and polygonal numbers

Abstract.
A linear combination $aT_{r}(m)+bT_{s}(n)$ of an $r$-gonal number $T_{r}(m)$ and an $s$-gonal number $T_{s}(n)$ with mutually coprime positive integer coefficients $a$ and $b$ produces infinitely many primes as $m$ and $n$ varies over the natural numbers, whereas the sum of the reciprocals of such primes converges unless $T_{r}(m)=m^{2}$ and $T_{s}(n)=n^{2}$. For each pair of coprime positive integers $a$ and $b$, there are arbitrary long arithmetic progressions among the primes of the form $am^{2}+bn^{2}$.
Key words and phrases:
prime, polygonal number, triangular number, arithmetic progression
2020 Mathematics Subject Classification:
Primary 11N32, 11N13, 11N36; Secondary 11P32, 11B25, 11B05
1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $4k+1$. Since there are infinitely many primes of the form $4k+1$, it follows that there are infinitely many primes which are sum of two squares. Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz. $n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions. However, it turns out not to be the case. In fact, the sum of the reciprocals of such primes converges. One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
$T_{r}(m)\;:=\;(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT_{r}(m)+bT_{s}(n).$ We answer this in the affirmative.
Theorem 1.
Given a pair of integers $r,s>2$ and a pair of mutually coprime positive integers $a$ and $b$, for all $m,n\in\mathbb{N}$, the number of primes of the form
$aT_{r}(m)+bT_{s}(n)$ (1.2)
which are less than $N$ is
$\asymp\begin{cases}\dfrac{N}{\log N},&\text{ if }r=s=4\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{otherwise.}\end{cases}$
Recall that $f\asymp g$ if and only if there exists a pair of constants $C,C^{\prime}>0$ such that $Cg\leq f\leq C^{\prime}g$. In the first case above, we have $T_{4}(m)=m^{2}$. The surprise is in the other case, where even if one of the summands is a multiple of a square, the corresponding set of primes has a lower rate of growth. From the above theorem, via Abel’s summation formula, we obtain
Corollary 1.
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT_{r}(m)+bT_{s}(n)$ diverges if and only if $r=s=4$.
In particular, Theorem 1 answers a question from [1], which asks how many primes are sums of two triangular numbers.
Corollary 2.
There are infinitely many primes which are sum of two triangular numbers.
In fact, there are infinitely many such primes in each of the residue classes $\pm 1$ modulo $4$.
Theorem 2.
Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes $\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
$\asymp\dfrac{N}{(\log N)^{3/2}}.$
In particular, for $a=b=1$ and for the primes congruent to $1$ modulo $4$, we have the following.
Corollary.
For each odd integer $r>1$, the number of primes $p\leq N$, which are sums of two $r$-gonal numbers and also sums of two squares, is
$\asymp\frac{N}{(\log N)^{3/2}}.$
Since a number $n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares*** It follows from (1.1) that for all $r$, we have $4(r-2)(T_{r}(m)+T_{r}(n))+(r-4)^{2}\;=\;\left((r-2)(m-n)\right)^{2}+\left((r-2)(m+n)-r+4\right)^{2}.$ Conversely, if $4(r-2)M\;+\;(r-4)^{2}\;=\;A^{2}\;+\;B^{2}$ such that $A\;\pm\;B\;\equiv\;r\pmod{2(r-2)}$, then $M$ is a sum of two $r$-gonal numbers. (see [1]) and since a prime is a sum of two squares if and only if it is of the form $4k+1$ (see [8]), it follows that
Corollary.
The number of primes $p\leq N$ such that both $p$ and $4p+1$ are sums of two squares is
$\asymp\dfrac{N}{(\log N)^{3/2}}.$
By a primitive arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms. Now, looking at Theorem 2, one may wonder whether every primitive arithmetic progression contains infinitely many primes which are sums of two triangular numbers. However, that does not hold in general! For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $4n+1$ is not a sum of two squares. So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers. This prompts us to refine our query as follows:
Question 1.
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, if there is no congruence obstruction, does every primitive arithmetic progression contain infinitely many primes which are of the form $aT_{r}(m)+bT_{s}(n)?$
Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
Corollary 3.
For each pair of integers $r,s>2$, at least one of which is not equal to $4$ and for every pair of $a,b\in\mathbb{N}$, the double sequence
$\{aT_{r}(m)+bT_{s}(n)\}_{m,n\in\mathbb{N}}$
misses infinitely many primes from every primitive arithmetic progression.
In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given primitive arithmetic progression.
Theorem 1 along with Chebyshev’s theorem (Corollary 2.6 in [7]) and Green-Tao Theorem [4] implies:
Corollary 4.
For each pair of coprime positive integers $a$ and $b$, there are arbitrariliy long arithmetic progressions among the primes of the form
$am^{2}+bn^{2}.$
It also follows from Theorem 1 and Chebyshev’s theorem that the sequence of the primes, which are linear combinations (with coprime positive integer coefficients) of two polygonal numbers at least one of which is not a square, has zero relative density with respect to the sequence of all primes. So, possible existence of arbitrary long arithmetic progressions among such primes (see the following table) is undetectable by Green-Tao Theorem [4].
Table 1. Examples of arithmetic progressions among the primes which are sums of two triangular numbers
Length
Arithmetic progressions
3 31, 37, 43
4 43, 61, 79, 97
5 61, 151, 241, 331, 421
6 7, 37, 67, 97, 127, 157
7 157, 37747, 75337, 112927, 150517, 188107, 225697
8 3307, 69457, 135607, 201757, 267907, 334057, 400207, 466357
9 3823, 6133, 8443, 10753, 13063, 15373, 17683, 19993, 22303
10
73091, 135461, 197831, 260201, 322571, 384941, 447311,
509681, 572051, 634421
The following table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number. More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
Question 2.
For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT_{r}(m)+bT_{s}(n)?$
2. Proofs of the theorems
The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below. A
Theorem A (Theorem 1 in [5]).
Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\gg\;\;\frac{N}{\log N},$ (2.1)
if $\Delta:=b^{2}-4ac$ is a perfect square or $D:=af^{2}-bef+ce^{2}+g\Delta=0$, and
$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\asymp\;\;\frac{N}{(\log N)^{3/2}},$ (2.2)
if $\Delta$ is not a perfect square and $D\neq 0$.
Lemma B (Lemma 1 in [5]).
B Let $P(x,y)\;=\;ax^{2}+bxy+cy^{2}+ex+fy+g$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
$\Delta\;=\;b^{2}-4ac\;=\;0,\ \alpha\;:=\;bf-2ce\;=\;0\text{ \ and \ }\beta\;:=\;be-2af\;=\;0.$
Proof of Theorem 1. For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
$\mu T_{r}(m)+\nu T_{s}(n),$
we note that (see (1.1)) the above linear combination is equal to
$P_{1}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu(4-s)y$ (2.3)
for m= 2xm\;=\;2x and n= 2yn\;=\;2y,
$P_{2}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu(4-r)x+\nu sy+\nu$ (2.4)
for m= 2xm\;=\;2x and n= 2y+1n\;=\;2y+1,
$P_{3}(x,y)\;:=\;2\mu(r-2)x^{2}+2\nu(s-2)y^{2}+\mu rx+\nu sy+\mu+\nu$ (2.5)
for m= 2x+1m\;=\;2x+1 and n= 2y+1n\;=\;2y+1.
Since the coefficients of $x^{2}$ and $y^{2}$ in each of the above polynomials are positive and since the coefficient of $xy$ is zero, it follows that
$\Delta\;<\;0$ (2.6)
for each of the above polynomials (see Lemma B). Hence, $\frac{\partial{P_{i}}}{\partial x}$ and $\frac{\partial{P_{i}}}{\partial y}$ linearly independent for all $i\in\{1,2,3\}$.
Also, from (2.3), (2.4) and (2.5), we note that for each of the polynomials $P_{1},P_{2}$ and $P_{3}$, we have
$D\;=\;2\mu\nu^{2}(r-2)(4-s)^{2}+2\mu^{2}\nu(s-2)(4-r)^{2},$ (2.7)
which equals zero if and only if $r=s=4$ (see Theorem A).
In order to apply Theorem A, we must also check whether the polynomials $P_{1},P_{2}$ $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$. Suppose, for some $i\in\{1,2,3\}$, there exists $a,b,c,a^{\prime},b^{\prime},c^{\prime}\in\mathbb{Z}$ such that
$P_{i}(x,y)\;=\;(ax+by+c)(a^{\prime}x+b^{\prime}y+c^{\prime}).$
Comparing the coefficients of $x^{2}$, $y^{2}$ and $xy$ in both of the above expressions for $P_{i}(x,y)$, we obtain
$aa^{\prime}\;=\;2\mu(r-2),\ bb^{\prime}\;=\;2\nu(s-2)\text{ and }ab^{\prime}+a^{\prime}b\;=\;0.$
Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$. Hence, substituting $kb$ for $a$ and $kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
$k^{2}\;=\;-\frac{\mu(r-2)}{\nu(s-2)}<0,$
which is absurd! Thus, $P_{i}(x,y)$ is irreducible in $\mathbb{Q}[x,y]$ for all $i\in\{1,2,3\}$. Hence, it follows from Gauss’s Lemma (see Ch.VI, §2 in [6]) that $P_{1},P_{2}$ and $P_{3}$ are irreducible in $\mathbb{Z}[x,y]$ if and only if these are primitive, i.e. the greatest common divisor of the coefficients of each of these polynomials is 1.
Since for all $i\in\{1,2,3\}$, the coefficients of $x^{2}$ and $y^{2}$ in $P_{i}(x,y)$ are positive and even, whereas the coefficient of $xy$ is zero, it follows that if any of the remaining coefficients of $P_{i}$ are odd, then $P_{i}(x,y)$ produces infinitely many odd numbers for $x,y\in\mathbb{N}$. In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers. In the following we compute the gcd $G$ of the coefficients of $P_{1}$, $P_{2}$ and $P_{3}$. For $i\in\{1,2,3\}$ and $N\in\mathbb{N}$, let
$\pi_{i}(N)\;:=\;\sum_{{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}}1.$
Case I. For $P_{1}$, from (2.3) we note that
$G\;=\;\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu(4-s))\;=\;\gcd(4,\mu r,\nu s).$
Hence, $G=1$ if and only if either $\mu r$ or $\nu s$ is odd. In particular, $G=1$ implies $D>0$ (see (2.7)). So, it follows from Theorem A that
$\pi_{1}(N)\;\asymp\;\begin{cases}\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
Case II. For $P_{2}$, from (2.4) we note that
$G\;=\;\gcd(2\mu(r-2),2\nu(s-2),\mu(4-r),\nu s,\nu)\;=\;\gcd(4,r,\nu).$
Hence, $G=1$ if and only if either $r$ or $\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\pi_{2}(N)\;\asymp\;\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
Case III. For $P_{3}$, from (2.5) we note that
$G\;=\;\gcd(2\mu(r-2),2\nu(s-2),\mu r,\nu s,\mu+\nu)\;=\;\gcd(4,r,s,\mu+\nu).$
Hence, $G=1$ if and only if either $r$ or $s$ of $\mu+\nu$ is odd. So, it follows from (2.7), Theorem A and Chebyshev’s theorem (Corollary 2.6 in [7]) that
$\pi_{3}(N)\;\asymp\;\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
Since either $\nu$ or $\mu+\nu$ is odd for all mutually coprime pairs $\mu$ and $\nu$, the claim follows from the above cases. ∎
[…]
Acknowledgments
The authors would like to thank Prof. Don Zagier for his comments and suggestions. All the computations for Table 1 was done using PARI/GP. Also, Bhattacharya thankfully acknowledges the support from the Department of Science and Technology, Govt. of India through the INSPIRE Faculty Award IFA17-MA104 and Rahaman thankfully acknowledges the support from the Ministry of Education, Govt. of India through the Prime Minister Research Fellowship 0501972.